# Privacy Beyond Pixels: Latent Anonymization for Privacy-Preserving Video Understanding

### v1

@@ 4.1 Privacy Evaluation
First, we employ an established privacy preservation protocol to ensure that $f_{A}$ removes sensitive-attribute related information. Although we focus on action-related video understanding models, private attribute information is still exposed in the latent features of the backbone encoder. The VISPR dataset orekondy2017towards evaluates privacy-preservation by measuring performance on a multi-class classification problem for various sensitive visual private attributes, with performance measured by mean average precision across classes (cMAP). Since the goal is to enhance privacy—not to accurately predict attributes—a lower performance indicates better privacy preservation. For evaluation, we train a linear classifier on static clip representations, adhering to protocols from prior work in video privacy wu2020privacy; dave2022spact. Additionally, we evaluate the effect of anonymization on a motion-based sensitive attribute, namely gait recognition on Casia-B yu2006framework.
@@ 4.2 Utility Video Task Evaluation
Action Recognition Action recognition involves analyzing spatio-temporal video features to classify actions. In our framework, features are extracted from all videos using a Kinetics-pretrained video encoder. Then, a linear classifier is trained for evaluation on each dataset $\mathbb{D}_{reco}$, namely Kinetics400 carreira2017quo, UCF101 soomro2012ucf101, and HMDB51 kuehne2011hmdb. Evaluation is top-1 accuracy on 5 evenly spaced clips from each test video.
Temporal Action Detection Temporal action detection (TAD) is a task that involves identifying the specific time intervals within an untrimmed video where particular actions occur. TAD utilizes features from a Kinetics-pretrained video encoder model. Given $\mathbb{D}_{tad}$, $f_{A}$ is used to generate anonymized feature set $\mathbb{F}_{tad}=\{\,f_{A}(f_{E}(X^{(i)}))\;|\;\forall X^{(i)}\in\mathbb{D}_{tad}\,\}$. Our TAD evaluation uses THUMOS14 THUMOS14 as $\mathbb{D}_{tad}$. We choose one of the recent state-of-the-art methods, TriDet shi2023tridet with default hyperparameters to evaluate using mean Average Precision (mAP).
Weakly-Supervised Anomaly Detection Weakly supervised anomaly detection (WSAD) involves localizing timestamps of anomalous (unexpected) events given long, untrimmed videos and only video-level labels. Our evaluation uses UCF-Crime sultani2018real as $\mathbb{D}_{anomaly}$. Given $\mathbb{D}_{anomaly}$, $f_{A}$ is used to create an anonymized feature set $\mathbb{F}_{anomaly}=\{\,f_{A}(f_{E}(X^{(i))}))\;|\;\forall X^{(i)}\in\mathbb{D}_{anomaly}\,\}$. A recent state-of-the-art anomaly detection method MGFN chen2023mgfn is used with default hyperparameters. Final evaluation is given as a frame-level ROC AUC percentage.
@@ 4.3 Gender Presentation Bias Protocols
NTU Bias Evaluation We further verify anonymization efficacy by evaluating on our proposed attribute bias protocols. The NTU60 shahroudy2016ntu action recognition dataset is curated to have minimal scene and subject biases as each actor performs each action in different scenes. Given that each video is labeled with subject ID, we can introduce an artificial bias by controlling the gendered subclass ratios across actions. A ratio of 95% sagawa2019distributionally is set for all but one action, where the typical ratio is inverted to create a spurious shortcut for the model. This is done for each gender, resulting in two subsets: NTU-Bias-F and NTU-Bias-M. Detailed protocol creation info is found in Appendix Sec. A.
@@
Anonymization
Method
Network Privacy
@@ 5.1 Main Evaluation: Privacy vs Task Tradeoffs
Our evaluation of the proposed method covers privacy protocols and a variety of downstream tasks. We observe in Table 1 that our approach consistently generalizes well across all tasks, closely maintaining the performance of the non-anonymized videos. In contrast, previous methods struggle to preserve performance uniformly across tasks, evident in the temporal action detection results of wu2020privacy; dave2022spact; fioresi2023ted. Experiments with large VFMs see similar performance trends, confirming the efficacy and scalability of SPLAVU.
@@ 5.2 Gender Bias Evaluation
The first row of Table 2 shows the performance difference between each gender presentation subclass in the NTU-Bias-F protocol, where the action brush_hair is chosen as the gendered shortcut action label. The baseline performance disparity between perceived gender subclasses is an unacceptably large 9.42%. Applying latent anonymization impressively reduces this gap by a relative 42.3%. The second row includes results for the complimentary protocol NTU-Bias-M (also brush_hair shortcut). Interestingly, the baseline subclass performance disparity is less than that of NTU-Bias-F (5.00%), but our method is still capable of reducing this unfair split and improving overall performance.
To confirm that these observations hold true in a real-world setting, we look at the final row of Table 2 to see the performance on the TSH Das_2019_ICCV protocol. Notably, our method improves the both the classifier quality and fairness. In this realistic scenario with a naturally occurring bias, SPLAVU reduces the gap between perceived gender subclasses by an astonishing relative 39.5%.
@@ 5.3 Ablations and Analysis
Pretraining
Dataset
VISPR K400 UCF101 HMDB51 ToyotaSH UCF-Crime THUM14
@@
Effect of task-specific training: Our important ablation in Table 4 demonstrates the effects of training our anonymizer without specific tasks. Notably, the highlighted cells show impressive generalization to unseen tasks with just a minor drop in performance compared to training on them. For example, looking at row (c) shows $f_{A}$ training with only action detection, yet the performance on action recognition and anomaly detection remain within 1.3% of the non-anonymized score. Across the board, thanks to the latent consistency loss, performance is not dependent on having seen the given utility task during training, proving that SPLAVU can effectively generalize to unseen tasks.
@@
Temporal sensitive attribute recognition: Table 6 shows the performance of our model on the retrieval-based gait recognition task with no training. Because gait recognition benefits from understanding a temporal signature, latent consistency does not suppress potentially sensitive temporal attributes. If task overfitting is not a concern, not using latent consistency properly defends against private temporal attribute recognition. Otherwise, a temporal sensitive attribute prediction task can be additionally included in the budget privacy loss wu2020privacy to maintain generalization.
@@
Limitations: The latent anonymization framework cannot mitigate a threat that may occur if videos must be transmitted before feature extraction. We leave such cases to complementary techniques like secure transmission, or to prior anonymization works if optimal accuracy is not a concern.
@@ Appendix A Dataset Details
Kinetics400 carreira2017quo is a large-scale video action dataset of YouTube videos which includes 400 human action classes with at least 400 video clips for each action. Each clip lasts around 10 seconds and is labeled with a single action class. The dataset is widely used for pretraining deep learning models for use in many video understanding tasks.
@@
UCF101 soomro2012ucf101 is an action recognition dataset of realistic action videos consisting of 101 action categories. With over 13,000 videos from various actions and scenes, it provides a diverse set of actions and a broad range of variability in terms of actions, viewpoints, appearances, and backgrounds.
@@
HMDB51 kuehne2011hmdb is a collection of 6,766 video clips distributed across 51 human action categories, each containing a minimum of 101 clips. The dataset includes a wide range of human actions and is designed for the development and evaluation of action recognition methods.
@@
NTU RGB+D 60 shahroudy2016ntu is a large-scale multi view human action recognition dataset complete with RGB video, depth maps, and skeleton joints, and IR sequences. This work only uses the RGB frames. Each of the 40 subjects are recorded completing 60 daily activities from 3 different cameras.
@@
Toyota Smarthome Das_2019_ICCV is a challenging real-world activity classification dataset captured from 7 independent Kinect v1 cameras. The clips recorded 18 senior subjects performing 31 daily activities in a natural manner. This work only uses the provided RGB frames. The dataset contains high class imbalance, intra-class variation, and duration variance.
@@
THUMOS14 THUMOS14 focuses on temporal action localization in untrimmed videos. It extends the UCF101 dataset with temporal annotations for a subset of the action classes, providing detailed temporal annotations for 20 action classes across 200 validation videos and 213 test videos.
@@
UCF-Crime sultani2018real is a large-scale dataset of surveillance videos designed for anomaly detection. It consists of 1,900 long and untrimmed videos for a total of 128 hours. The videos contain examples of 13 different real-world anomalies, including burglary, robbery, and fighting, among others, making it suitable for training and evaluating video anomaly detection models.
@@
VISPR orekondy2017towards consists of around 22,000 Flickr images annotated with 68 privacy-related attributes such as gender, age group, skin color, and more. It offers a multi-class classification protocol for assessing private attribute prediction. Table 7 shows the VISPR attribute split used, which we have adopted from wu2020privacy; dave2022spact; fioresi2023ted.
@@
Casia-B yu2006framework is a gait recognition dataset consisting of 13,640 video clips of 124 subjects. We utilize it to evaluate the effect of anonymization on a temporal-based sensitive attribute, namely gait. Evaluation is through top-1 retrieval of a gallery video with the same subject. The full dataset is used for evaluation, with the first 4 repetitions of each walk in the gallery and the last 2 per subject used as the probe.
@@
Table 7: Privacy attributes from subset of VISPR orekondy2017towards labels as used in previous works.
VISPR1 wu2020privacy; dave2022spact; fioresi2023ted
@@
Proposed NTU Bias Evaluation Details More formal details for the creation of the proposed perceived gender NTU bias protocol are described here. While the original dataset is balanced in terms of scene and actor, the distribution of actor/video counts are not balanced with respect to perceived gender. To properly evaluate bias mitigation, it is essential to ensure that there are no performance differences stemming from the larger number of male subjects and training videos. The subject IDs are used to first restructure the dataset in an effort to maximize fairness across the gender subgroups. As such, within themselves, the train and test sets should contain both an even number of male and female subjects AND an even number of videos per action. Formally, lets take the set of subjects $S=\{s_{i}\}^{N_{S}}_{i=1}$, where $N_{S}$ is the number of subjects in the dataset. For each subject $s_{i}\in S$, there is an associated gender label $\mathbf{g}(s_{i})$ where $\mathbf{g}(s_{i})\in\{male,female\}$. We set $N_{m}=N_{f}=\frac{N_{S}}{2}$, where $N_{m}$ and $N_{f}$ are the number of male and female subjects, respectively. Using the above notation with $\mathbb{D}_{IID}$ abbreviated to $D$, we define $D_{m}=\{(\mathbf{x}_{i},\mathbf{y}_{i},\mathbf{g}_{i})\in D|\mathbf{g}_{i}=male\}$ and $D_{f}=\{(\mathbf{x}_{i},\mathbf{y}_{i},\mathbf{g}_{i})\in D|\mathbf{g}_{i}=female\}$. We set $|D_{m}|=|D_{f}|=\frac{|D|}{2}$. With the dataset balanced across subject counts, subject genders, video count per action/gender, and background representation, the model should not have access to simple bias shortcuts.
To directly measure gender presentation bias, we inject an artificial bias related to perceived gender by creating a simple spurious shortcut for the model to follow. Specifically, we control the subclass ratios across all actions, setting $P(\mathbf{g}(s)=male|\mathbf{y})=0.95$ and $P(\mathbf{g}(s)=female|\mathbf{y})=0.05$, following the correlation strength in sagawa2019distributionally. However, for one action chosen at random, we flip this ratio, keeping 95% of perceived female videos ($P(\mathbf{g}(s)=female|\mathbf{y})=0.95$) and only 5% of perceived male videos ($P(\mathbf{g}(s)=male|\mathbf{y})=0.05$). We refer to this subset as NTU-Bias-F. To ensure that the shortcut taking is gender presentation agnostic, we repeat this protocol by swapping the subclasses, creating NTU-Bias-M. We find that swapping this subclass ratio for one action class reduces overall performance and causes a gap in subclass performance.

### v2

@@ 4.1 Privacy Evaluation
First, we employ an established privacy preservation protocol to ensure that $f_{A}$ removes sensitive-attribute related information. Although we focus on action-related video understanding models, private attribute information is still exposed in the latent features of the backbone encoder. The VISPR dataset Orekondy et al. (2017) evaluates privacy-preservation by measuring performance on a multi-class classification problem for various sensitive visual private attributes, with performance measured by mean average precision across classes (cMAP). For evaluation, we train a classifier on static clip representations, adhering to protocols from prior work in video privacy Dave et al. (2022b). Additionally, we evaluate on temporal-based VP-HMDB51 and VP-UCF101 datasets Li et al. (2023b), which contain both action and private attribute labels. Here, the full clip representations are used for attribute classification with cMAP as the evaluation metric. Since the goal is to enhance privacy, not to accurately predict attributes, a lower performance indicates better privacy preservation. For further temporal analysis, we evaluate the effect of anonymization on a motion-based sensitive attribute, namely gait recognition on Casia-B Yu et al. (2006).
@@ 4.2 Utility Video Task Evaluation
Action Recognition Action recognition involves analyzing spatio-temporal video features to classify actions. In our framework, features are extracted from all videos using a Kinetics-pretrained video encoder. Then, a linear classifier is trained for evaluation on each dataset $\mathbb{D}_{AR}$, namely Kinetics400 Carreira and Zisserman (2017), UCF101 Soomro et al. (2012), and HMDB51 Kuehne et al. (2011). Additional action results are shown on VP-HMDB51 and VP-UCF101 Li et al. (2023b). Evaluation is top-1 accuracy on 5 evenly spaced clips from each test video.
Temporal Action Detection Temporal action detection (TAD) involves identifying the specific time intervals within an untrimmed video where particular actions occur. TAD utilizes features from a Kinetics-pretrained video encoder model. Given $\mathbb{D}_{TAD}$, $f_{A}$ is used to generate anonymized feature set $\mathbb{F}_{TAD}=\{\,f_{A}(f_{E}(X^{(i)}))\;|\;\forall X^{(i)}\in\mathbb{D}_{TAD}\,\}$. Our TAD evaluation uses THUMOS14 Jiang et al. (2014) as $\mathbb{D}_{TAD}$. We choose one of the recent state-of-the-art methods, TriDet Shi et al. (2023) with default hyperparameters to evaluate using mean Average Precision (mAP).
Weakly-Supervised Anomaly Detection Weakly supervised anomaly detection (WSAD) involves localizing timestamps of anomalous (unexpected) events given long, untrimmed videos and only video-level labels. Our evaluation uses UCF-Crime Sultani et al. (2018) as $\mathbb{D}_{AD}$. Given $\mathbb{D}_{AD}$, $f_{A}$ is used to create an anonymized feature set $\mathbb{F}_{AD}=\{\,f_{A}(f_{E}(X^{(i))}))\;|\;\forall X^{(i)}\in\mathbb{D}_{AD}\,\}$. A recent state-of-the-art anomaly detection method MGFN Chen et al. (2023) is used with default hyperparameters. Final evaluation is given as a frame-level ROC AUC percentage.
@@ 4.3 Gender Presentation Bias Protocols
NTU Bias Evaluation We further verify anonymization efficacy by evaluating on our proposed attribute bias protocols. The NTU60 Shahroudy et al. (2016) action recognition dataset is curated to have minimal scene and subject biases as each actor performs each action in different scenes. Given that each video is labeled with subject ID, we can introduce an artificial bias by controlling the gendered subclass ratios across actions. A ratio of 95% Sagawa et al. (2019) is set for all but one action, where the typical ratio is inverted to create a spurious shortcut for the model. This is done for each gender, resulting in two subsets: NTU-Bias-F and NTU-Bias-M. Detailed protocol creation info is found in Appendix Sec. A.
@@
Anonymization Method Network Privacy
@@
Table 2: Accuracy and privacy evaluation using video-level private attribute protocols VP-HMDB51 and VP-UCF101 Li et al. (2023b), evaluated as top-1 accuracy and cMAP. Our anonymizer successfully mitigates temporal private attribute prediction, matching the quality of a latent anonymizer trained with a supervised privacy loss Wu et al. (2020).
Anonymization Method VP-HMDB51 VP-UCF101
[…]
Raw Videos 72.6 76.4 96.8 75.9
Supervised 72.4 70.4 96.5 69.5
Ours 72.1 70.5 96.8 69.6
@@ 5.1 Main Evaluation: Privacy vs Task Tradeoffs
Our evaluation of the proposed method covers private attribute prediction protocols and a variety of downstream tasks. We observe in Table 1 that our approach consistently generalizes well across all tasks, closely maintaining the performance of the non-anonymized videos. In contrast, previous methods struggle to preserve performance uniformly across tasks, evident in the temporal action detection results of Wu et al. (2020); Dave et al. (2022b); Fioresi et al. (2023). Table 2 demonstrates resistance to temporal private attributes and that performance matches that of a version trained with supervised attribute labels. Experiments with large VFMs see similar performance trends, confirming the efficacy and scalability of SPLAVU. Results in Table 2 demonstrate the ability of our method to handle temporal private attribute settings.
@@ 5.2 Gender Bias Evaluation
The first row of Table 3 shows the performance difference between each gender presentation subclass in the NTU-Bias-F protocol, where the action brush_hair is chosen as the gendered shortcut action label. The baseline performance disparity between perceived gender subclasses is an unacceptably large 9.42%. Applying latent anonymization impressively reduces this gap by a relative 42.3%. The second row includes results for the complementary protocol NTU-Bias-M (also brush_hair shortcut). Interestingly, the baseline subclass performance disparity is less than that of NTU-Bias-F (5.00%), but our method is still capable of reducing this unfair split and improving overall performance.
To confirm that these observations hold true in a real-world setting, we look at the final row of Table 3 to see the performance on the TSH Das et al. (2019) protocol. Notably, our method improves the both the classifier quality and fairness. In this realistic scenario with a naturally occurring bias, SPLAVU reduces the gap between perceived gender subclasses by an astonishing relative 39.5%.
@@ 5.3 Ablations and Analysis
Pretraining Dataset VISPR K400 UCF101 HMDB51 ToyotaSH UCF-Crime THUM14
@@
Effect of task-specific training: Our important ablation in Table 5 demonstrates the effects of training our anonymizer without specific tasks. Notably, the highlighted cells show impressive generalization to unseen tasks with just a minor drop in performance compared to training on them. For example, looking at row (c) shows $f_{A}$ training with only action detection, yet the performance on action recognition and anomaly detection remain within 1.3% of the non-anonymized score. Across the board, thanks to the latent consistency loss, performance is not dependent on having seen the given utility task during training, proving that SPLAVU can effectively generalize to unseen tasks. See Appendix Table 13 for additional generalization results on the video retrieval task.
@@
Temporal sensitive attribute recognition: Table 7 shows the performance of our model on the retrieval-based gait recognition task with no training. Because gait recognition benefits from understanding a temporal signature, latent consistency does not suppress potentially sensitive temporal attributes. If task overfitting is not a concern, not using latent consistency properly defends against private temporal attribute recognition. Otherwise, a temporal sensitive attribute prediction task can be additionally included in the budget privacy loss Wu et al. (2020) to maintain generalization.
@@
Table 8: Feature inversion attack results. Our model demonstrates strong robustness to black-box attacks, where action performance is recovered but not private attribute prediction. Random chance on each dataset is included to contextualize performance.
Method/Attack VP-HMDB51 VP-UCF101
[…]
Raw Videos 72.6 76.4 96.8 75.9
Random Chance 2.0 58.3 1.0 58.7
Ours 72.1 70.5 96.8 69.6
Black-Box Inversion 67.3 58.8 96.8 58.4
White-Box Inversion 68.0 73.7 94.5 72.6
Attack resistance: A major concern of any anonymization framework is the potential for an inversion attack. Table 8 includes results from white and black-box attack settings on VP-HMDB51 and VP-UCF101 Li et al. (2023b). In the black-box setting, the attacker observes input-output pairs of the anonymizer and trains a separate inversion model where no gradients through $f_{A}$ are available. In the white-box setting, the attacker has full access to $f_{A}$ and instead performs gradient-based reconstruction by optimizing a synthetic embedding to match the observed anonymized output. In both cases, recovered embeddings are evaluated by feeding them into downstream models trained on non-anonymized features to measure action accuracy and privacy cMAP. Notably, in the black-box setting, the inversion network successfully recovers action classification performance, yet does not improve attribute classification performance above random chance. The model is not robust to white box attacks, which are typically unrealistic in deployment scenarios. Robustness to white-box attacks is a promising avenue for future work. See Appendix Sec. C and Figure 4 for qualitative analysis on reconstruction-based attacks, where our anonymization prevents faithful video reconstruction.
Limitations: The latent anonymization framework cannot mitigate a threat that may occur if videos must be transmitted before feature extraction. We leave such cases to complementary techniques like secure transmission, or to prior anonymization works if optimal accuracy is not a concern. Our formulation currently targets global-embedding tasks (e.g., action recognition, temporal action detection, anomaly detection, retrieval). Extending anonymization to dense tasks such as segmentation or captioning that require patch-wise embeddings is an important direction for future work.
@@ Appendix A Dataset Details
Kinetics400 Carreira and Zisserman (2017) is a large-scale video action dataset of YouTube videos which includes 400 human action classes with at least 400 video clips for each action. Each clip lasts around 10 seconds and is labeled with a single action class. The dataset is widely used for pretraining deep learning models for use in many video understanding tasks.
@@
UCF101 Soomro et al. (2012) is an action recognition dataset of realistic action videos consisting of 101 action categories. With over 13,000 videos from various actions and scenes, it provides a diverse set of actions and a broad range of variability in terms of actions, viewpoints, appearances, and backgrounds.
@@
HMDB51 Kuehne et al. (2011) is a collection of 6,766 video clips distributed across 51 human action categories, each containing a minimum of 101 clips. The dataset includes a wide range of human actions and is designed for the development and evaluation of action recognition methods.
@@
NTU RGB+D 60 Shahroudy et al. (2016) is a large-scale multi view human action recognition dataset complete with RGB video, depth maps, and skeleton joints, and IR sequences. This work only uses the RGB frames. Each of the 40 subjects are recorded completing 60 daily activities from 3 different cameras.
@@
Toyota Smarthome Das et al. (2019) is a challenging real-world activity classification dataset captured from 7 independent Kinect v1 cameras. The clips recorded 18 senior subjects performing 31 daily activities in a natural manner. This work only uses the provided RGB frames. The dataset contains high class imbalance, intra-class variation, and duration variance.
@@
THUMOS14 Jiang et al. (2014) focuses on temporal action localization in untrimmed videos. It extends the UCF101 dataset with temporal annotations for a subset of the action classes, providing detailed temporal annotations for 20 action classes across 200 validation videos and 213 test videos.
@@
UCF-Crime Sultani et al. (2018) is a large-scale dataset of surveillance videos designed for anomaly detection. It consists of 1,900 long and untrimmed videos for a total of 128 hours. The videos contain examples of 13 different real-world anomalies, including burglary, robbery, and fighting, among others, making it suitable for training and evaluating video anomaly detection models.
@@
VISPR Orekondy et al. (2017) consists of around 22,000 Flickr images annotated with 68 privacy-related attributes such as gender, age group, skin color, and more. It offers a multi-class classification protocol for assessing private attribute prediction. Table 9 shows the VISPR attribute split used, which we have adopted from Wu et al. (2020); Dave et al. (2022b); Fioresi et al. (2023).
@@
Casia-B Yu et al. (2006) is a gait recognition dataset consisting of 13,640 video clips of 124 subjects. We utilize it to evaluate the effect of anonymization on a temporal-based sensitive attribute, namely gait. Evaluation is through top-1 retrieval of a gallery video with the same subject. The full dataset is used for evaluation, with the first 4 repetitions of each walk in the gallery and the last 2 per subject used as the probe.
VP-HMDB51/VP-UCF101 Li et al. (2023b) are versions of the base HMDB51 Kuehne et al. (2011) and UCF101 Soomro et al. (2012) datasets annotated with five clip-based visual private attribute labels, following the VISPR label setting.
@@
Table 9: Privacy attributes from subset of VISPR Orekondy et al. (2017) labels as used in previous works.
VISPR1 Wu et al. (2020); Dave et al. (2022b); Fioresi et al. (2023)
@@
Proposed NTU Bias Evaluation Details More formal details for the creation of the proposed perceived gender NTU bias protocol are described here. While the original dataset is balanced in terms of scene and actor, the distribution of actor/video counts are not balanced with respect to perceived gender. To properly evaluate bias mitigation, it is essential to ensure that there are no performance differences stemming from the larger number of male subjects and training videos. The subject IDs are used to first restructure the dataset in an effort to maximize fairness across the gender subgroups. As such, within themselves, the train and test sets should contain both an even number of male and female subjects AND an even number of videos per action. Formally, let’s take the set of subjects $S=\{s_{i}\}^{N_{S}}_{i=1}$, where $N_{S}$ is the number of subjects in the dataset. For each subject $s_{i}\in S$, there is an associated gender label $\mathbf{g}(s_{i})$ where $\mathbf{g}(s_{i})\in\{male,female\}$. We set $N_{m}=N_{f}=\frac{N_{S}}{2}$, where $N_{m}$ and $N_{f}$ are the number of male and female subjects, respectively. Using the above notation with $\mathbb{D}_{IID}$ abbreviated to $D$, we define $D_{m}=\{(\mathbf{x}_{i},\mathbf{y}_{i},\mathbf{g}_{i})\in D|\mathbf{g}_{i}=male\}$ and $D_{f}=\{(\mathbf{x}_{i},\mathbf{y}_{i},\mathbf{g}_{i})\in D|\mathbf{g}_{i}=female\}$. We set $|D_{m}|=|D_{f}|=\frac{|D|}{2}$. With the dataset balanced across subject counts, subject genders, video count per action/gender, and background representation, the model should not have access to simple bias shortcuts.
To directly measure gender presentation bias, we inject an artificial bias related to perceived gender by creating a simple spurious shortcut for the model to follow. Specifically, we control the subclass ratios across all actions, setting $P(\mathbf{g}(s)=male|\mathbf{y})=0.95$ and $P(\mathbf{g}(s)=female|\mathbf{y})=0.05$, following the correlation strength in Sagawa et al. (2019). However, for one action chosen at random, we flip this ratio, keeping 95% of perceived female videos ($P(\mathbf{g}(s)=female|\mathbf{y})=0.95$) and only 5% of perceived male videos ($P(\mathbf{g}(s)=male|\mathbf{y})=0.05$). We refer to this subset as NTU-Bias-F. To ensure that the shortcut taking is gender presentation agnostic, we repeat this protocol by swapping the subclasses, creating NTU-Bias-M. We find that swapping this subclass ratio for one action class reduces overall performance and causes a gap in subclass performance.